### v4

# Jina-ColBERT-v2: A General-Purpose Multilingual Late Interaction Retriever

Abstract
Multi-vector dense models, such as ColBERT, have proven highly effective in information retrieval. ColBERT’s late interaction scoring approximates the joint query-document attention seen in cross-encoders while maintaining inference efficiency closer to traditional dense retrieval models, thanks to its bi-encoder architecture and recent optimizations in indexing and search. In this work we propose a number of incremental improvements to the ColBERT model architecture and training pipeline, using methods shown to work in the more mature single-vector embedding model training paradigm, particularly those that apply to heterogeneous multilingual data or boost efficiency with little tradeoff. Our new model, Jina-ColBERT-v2, demonstrates strong performance across a range of English and multilingual retrieval tasks.
1 Introduction
Neural retrieval has gained popularity in recent years following the arrival of capable pre-trained language models (PLMs) (Devlin et al., 2019; Liu et al., 2019; Clark et al., 2020). Two types of approaches have been employed to apply PLMs to retrieval. Sparse neural retrieval systems, such as SPLADE (Formal et al., 2021), represent texts as weighted bags of words that are interpreted as sparse high-dimensional vectors for maximum inner product search (MIPS). Dense retrievers similarly encode queries and documents as dense vectors, capturing relevance signals through spatial relationships extending beyond exact term matching.
Most dense retrievers encode a query or document as a single vector, commonly the result of mean-pooling or the [CLS]-embedding over the transformer’s final layer token embeddings. In contrast, recent multi-vector retrievers like ColBERT (Khattab and Zaharia, 2020) generalize this embedding process to maintain an embedding for each token, computing relevance scores as a function of the similarities of query and document tokens instead. To make the ColBERT usable in practice, the output dimensionality is restricted to be much smaller than the single-vector models. This approach has the benefit of remaining compatible with much of the vector similarity infrastructure that makes single-vector methods efficient, but requires more space to store even a smaller embedding per token and compute at inference time to aggregate token interactions into a single score. This late interaction over token embeddings achieves greater in-domain performance and tends to be more robust out-of-domain than single-vector similarity. While ColBERTv2 is trained only on English MSMARCO triplets (Bajaj et al., 2016) and has a monolingual BERT backbone, making it incapable of multilingual retrieval, some previous works extend the model to multilingual retrieval.
ColBERT-XM (Louis et al., 2024) does this by using parameter extensions for each additional language, and (Lawrie et al., 2023) trains solely on machine-translated English MSMARCO data to get effective heterogeneous multilingual performance. These approaches, however, come with trade-offs in terms of model usability and training data diversity. Other multilingual multi-vector models like BGE-M3 (Chen et al., 2024) produce extremely large token representations that limit their practical utility for first-stage retrieval.
In this work, we propose Jina-ColBERT-v2, which introduces an improved training recipe for ColBERT models with the following features:
Training with diverse weakly-supervised data: We additionally pretrain our modified PLM with rotary position embedding and train on large-scale unlabeled text pairs from various corpus with a weakly-supervised single-vector contrastive objective. A second-stage of ColBERT finetuning with labeled triplet data and supervised distillation is used to further boost its performance.
General multilingual performance: We train with data from a variety of high- and low-resource languages using both labeled and unlabeled data, including human- and machine-translated training data, and show that this improves even out-of-domain multilingual performance.
Inference-agnostic efficiency: We introduce multiple sizes of linear projection heads, jointly trained using the non-weight tying variant of Matryoshka Representation Loss Kusupati et al. (2022), enabling the selection of token embedding size at inference time with minimal performance degradation. We demonstrate that reducing the embedding dimensionality in half from 128 to 64 yields only a minor performance tradeoff. Additionally, our flash-attention optimized backbone, Jina-XLM-RoBERTa  provides further free performance improvement during inference.
Our experimental results show competitive retrieval performance across both English and multilingual benchmarks. We also present controlled experiments demonstrating the benefits, or lack thereof, of the training modifications we consider in developing our training recipe.
2 Related Work
In this section, we discuss related work in single- and multi-vector retrieval, as well as the non-English late-interaction retrievers from which our training recipe draws inspiration.
2.1 Single-Vector Retrieval
Single-vector encoder models have demonstrated their potential as general-purpose embedding models across a number of downstream tasks (Muennighoff et al., 2023). When used in a bi-encoder retrieval model, they asymmetrically encode queries and documents as separate dense vectors, and measure their pairwise relevance as the cosine similarity between the vectors. Owing to their strong in-domain performance and straightforward inference scheme, there has been a growing focus on improving their training. Studies demonstrate that large-scale unsupervised pair training utilizing in-batch negatives, followed by a small-scale triplet finetuning stage, significantly improves performance compared to a dense retriever trained solely on triplet data (Li et al., 2023; Günther et al., 2023). Other works have incorporated asymmetric task-specific instructions for queries and documents to further enhance performance (Wang et al., 2024) and demonstrated the efficacy of using synthetically generated training data, including using diverse task instructions and machine translations, to further improve model representations. (Wang et al., 2023; Lee et al., 2024)
2.2 Multi-Vector Retrieval
Multi-vector retrievers like ColBERT also employ a bi-encoder structure, but queries and passages are represented by a collection of smaller token embeddings rather than one large vector. As such, ColBERTv2’s training uses many of the same techniques as state-of-the-art single-vector models: cross-encoder distillation, multiple negatives per query, and self-mined hard negatives. Recent models have continued to improve on this training recipe, particularly for multilingual or non-English training. BGE-M3 (Chen et al., 2024) adopts the two-stage pairs-to-triplets training pipeline, and does self-knowledge distillation, treating the combination of its sparse, dense, and multi-vector scores as the teacher score.
2.3 Multilingual Retrieval
Owing to the quality of English-based pre-trained models (BERT) and annotated data (MSMARCO), many advances in neural retrieval have been applied first to the monolingual English setting Karpukhin et al. (2020); Xiong et al. (2020); Khattab and Zaharia (2020). Researchers, however, have also made advances in non-English capabilities.
On the modeling front, multilingual PLMs like mBERT (Devlin et al., 2019) and later XLM-RoBERTa (Conneau et al., 2020) have expanded pre-training to include text in up to 100 languages, including in cross-language contexts. For multilingual retrieval data, there are two approaches: natural and translated. Datasets like Mr-Tydi and MIRACL (Zhang et al., 2021, 2023b) are built from human-generated and annotated queries, whereas mMARCO (Bonifacio et al., 2022) is a collection of machine-translated copies of MSMARCO which inherit their judgments from the original dataset. The former method tends to be of higher quality and lacks the subtle distributional/idiomatic errors, dubbed "translationese", that the latter sometimes exhibits. Naturally, however, human generation costs more per example.
Recent multi-vector work has also proposed further modifications along the dimensions of architecture and data. ColBERT-XM Louis et al. (2024) addresses the so-called curse of multilinguality (Conneau et al., 2020), the performance degradation of models pre-trained on too many tasks, with shared- and per-language parameters that allow for more robust zero-shot language transfer and post-hoc language extension. On the data approach, ColBERT-X (Nair et al., 2022; Lawrie et al., 2023; Yang et al., 2024) uses language-mixed batches of machine-translated English data, and BGE-M3 (Chen et al., 2024) curates unsupervised and high-quality supervised corpora of diverse multilingual training data.
3 Training Overview
Jina-ColBERT-v2’s training paradigm has three parts:
  1. 1.
    Modified Encoder Architecture: We use a modified encoder backbone, derived from XLM-RoBERTa with improvements made to its architecture and pre-training regime. We further extend ColBERT’s linear projection head by jointly training a collection of different-size heads for embedding size reduction.
  2. 2.
    Pair Training: To learn from the semantic structure of large quantities of diverse data in many languages, we first train our encoder model on weakly supervised text pairs from a variety of embedding datasets.
  3. 3.
    Triplet Training: Our model is further finetuned using retrieval examples in many languages with both positives and hard negatives, supervised by a highly-capable multilingual cross-encoder.
The following sections describe our experiments on these three components of training Jina-ColBERT-v2.
4 Architecture
4.1 Backbone Improvements
Following many prior single- and multi-vector multilingual training efforts, we adopted XLM-RoBERTa as our backbone model due to its strong performance across various downstream tasks Nair et al. (2022); Louis et al. (2024); Chen et al. (2024). To improve the efficiency, we enhance the XLM-RoBERTa architecture with flash attention (Dao, 2024).
We replace the absolute positional embeddings with rotary positional embeddings (RoPE, Su et al. (2023)), which are empirically understood to be better. They also have the advantage of supporting context lengths far longer than 512 tokens, although we do not explicitly focus on long-context in this work. To warm up its new positional embeddings, we continued pre-training the modified backbone with the same masked language modeling objective for 160,000 steps on the RefinedWeb dataset Penedo et al. (2023), a modern, high-quality corpus, under the masked language modeling objective. During this pre-training phase, we set the maximum sequence length to 8,192 tokens with a rotary base of 10,000 and employed whole-word-masking Devlin et al. (2019), masking out 30% of the tokens. We call this modified language model Jina-XLM-RoBERTa.
4.2 Multiple Linear Heads
To reduce index sizes, ColBERT includes a linear head that projects its token embeddings from the hidden dimension of its language model down to a lower dimension ($768\to 128$). As a notable exception, BGE-M3’s multi-vector retrieval does not take this step, keeping its token embeddings at a full 1024 dimensions.
We jointly train six linear heads with dimensions $d\in\{64,96,128,256,512,768\}$ using Matryoshka Representation Loss (MRL, Kusupati et al. (2022)). This allows users to choose greater or lesser space efficiency, with an associated performance trade-off. Figure 1 quantifies this tradeoff, showing the strong performance preservation of our reduced-dimension linear heads. Halving the token dimension ($128\to 64$) only causes its nDCG@10 to drop by 0.01 (1.59%). We unfortunately find that MRL’s weight-tying efficient variant (MRL-E), where losses are computed on truncations of the same token vector does not preserve performance well, which we hypothesize is a consequence of the already-low projected dimension of the original ColBERT formulation.
Figure 1: nDCG@10 scores for BEIR datasets using 64-, 96-, and 128-dimension linear projection heads for token embeddings.
5 Pair Training
To leverage an abundance of text pairs with varying richness of semantic structure, we draw inspiration from common practices in single-vector embedding model training and begin with training on these text pairs, focusing on optimizing the embedding model’s performance on general semantic similarity and relatedness tasks. This weakly-supervised stage is in contrast to previous ColBERT works, which typically start directly from a PLM like BERT with triplet training on 32-way or 64-way retrieval triplets consisting of a query, a positive passage, and multiple mined negatives.
5.1 Data Composition
Our pair training data consists of a broad range of weakly supervised datasets harvested from the web. We adjusted sampling rates across different languages and domains based on intuition, resulting in a set of 450 million weakly supervised, semantically related sentence pairs, question-answer pairs, and query-document pairs. Of these 450 million pairs, 50.0% are in English. Our non-English pair-wise datasets contain a diverse collection of 29 major languages, including 3.0% code data, with 4.3% representing cross-lingual data.
5.2 Contrastive Loss
We utilize the same single-vector pair-training loss function as described in Günther et al. (2023). Due to the often symmetric nature of our text pairs, the loss is calculated in both directions. During the pair training stage, we set the temperature $\tau=0.02$ and used a peak learning rate of $5\times 10^{-5}$ with a warm-up period of 1,000 steps. The model was trained using the Adam optimizer for 100,000 steps with a global batch size of 16,384.
6 Triplet Training
6.1 Data Composition
Our triplet dataset consists of 1) high-quality, human-annotated research datasets such as MSMARCO, DuReader, and MIRACL (Bajaj et al., 2016; He et al., 2018; Zhang et al., 2023b) with diversely mined negatives 2) high-quality datasets like MSMARCO and NQ translated from English into Chinese, French, German, Japanese, Russian and Spanish, following our previous work Mohr et al. (2024) and 3) synthetically generated datasets to address common failure modes of dense vector models such as negation and to cover niche domains like legal IR.
The triplet dataset covers 14 widely used languages, with a strong emphasis on Arabic, Chinese, English, French, German, Japanese, Russian, and Spanish. We sample the datasets to create a language distribution similar to that used in pair training. English accounts for 45.9% of the triplets, with 52.1% roughly evenly split between the mentioned high-resource non-English languages and a small 2.0% share for lower-resource languages.
Notably, owing to the limitations of our various sources of data, we train on triplets with only 7 negatives per example, in contrast to the 32- or 64-way triplets of ColBERTv2.
6.2 Supervision Loss
Following ColBERTv2, we finetune our pair-trained checkpoint on samples with hard negatives using a KL divergence loss function to distill soft labels from the teacher model. For the teacher model, we use jina-reranker-v2-base-multilingual [^1] https://huggingface.co/jinaai/jina-reranker-v2-base-multilingual, a highly capable multilingual cross encoder.
This stage trains for 100,000 steps with a batch size of 32 and a cosine decay learning rate schedule with 5% warm-up that peaks at $1\times 10^{-5}$. We use pure BFLOAT-16 precision, and apply magnitude-based gradient clipping with a threshold of $1$ for stability.
7 Results
BEIR avg nf fi tc ar qu sd sf to db fe cf hp nq
BM25 44.0 32.5 23.6 65.6 31.5 78.9 15.8 66.5 36.7 31.3 75.3 21.3 60.3 32.9
ColBERTv2 49.6 33.7 35.4 72.6 46.5 85.5 15.4 68.9 26.0 45.2 78.5 17.6 67.5 52.4
answerai-v1 55.7 37.3 41.2 84.6 50.1 87.7 18.4 74.8 25.7 45.6 91.0 33.1 76.1 59.1
Ours 53.1 34.6 40.8 83.4 36.6 88.7 18.6 67.8 27.4 47.1 80.5 23.9 76.6 64.0
Table 1: Comparison of nDCG@10 scores between BM25, ColBERTv2, answer-colbert-small and Jina-ColBERT-v1 and Jina-ColBERT-v2 on the BEIR test set. nf for NFCorpus, fi for FIQA (Fact In Question Answering), tc for TREC-COVID (Text Retrieval Conference COVID), ar for Arguana, qu for Quora, sd for SciDocs, sf for SciFact, to for Webis-Touche, db for DBpedia-Entity, fe for FEVER (Fact Extraction and Verification), cf for Climate-FEVER, hp for HotpotQA, and nq for Natural Questions
We evaluate Jina-ColBERT-v2 on four widely used benchmarks, BEIR, LoTTE, and MIRACL and mMARCO. For general English performance, we use the same subset of 14 retrieval and text-similarity tasks from the BEIR benchmark as in Santhanam et al. (2022). Additionally, we assess performance on the LoTTE benchmark, which focuses on long-tail queries, and the MIRACL and mMARCO benchmarks Zhang et al. (2023b); Bonifacio et al. (2022), which assess non-English retrieval performance. We report nDCG@10 for the BEIR and MIRACL collections, MRR@10 for mMARCO, and Success@5 for LoTTE. Scores are reported on the test split for BEIR, development split for MIRACL and mMARCO, and search test split for LoTTE. We use the same maximum query/document lengths as reported in Santhanam et al. (2022), and use the default (32/300) for MIRACL and mMARCO.
LoTTE avg Life. Rec. Wri. Sci. Tech.
BM25 67.8 80.2 68.5 74.7 53.6 61.9
ColBERTv2 72.0 84.7 72.3 80.1 56.7 66.1
Ours 76.4 87.0 77.6 83.8 60.5 73.0
Table 2: Comparison of Success@5 of various models across different LoTTE search query subsets.
Table 1 shows Jina-ColBERT-v2’s strong English performance compared to ColBERTv2, while still trailing the monolingual answerai-colbert-small-v1. Notably, however, we perform well below ColBERTv2 on ArguAna (ar), which we might attribute to either its unusual task: counterargument retrieval being at odds with our retrieval-heavy triplet training data distribution, or as an indication of the limitation of our stronger augmentation attention (discussed in Section 8.4) when applied to much longer (300 token) queries. Similarly for LoTTE, we see in Table 2 an improvement over ColBERTv2.
MIRACL avg ar bn de es en fa fi fr hi id ja ko ru sw te th yo zh
BM25 38.5 48.1 50.8 22.6 31.9 35.1 33.3 55.1 18.3 45.8 44.9 36.9 41.9 33.4 38.3 49.4 48.4 40.6 18.0
mDPR-ZS 41.8 49.9 44.3 49.0 47.8 39.4 48.0 47.2 43.5 38.3 27.2 43.9 41.9 40.7 29.9 35.6 35.8 39.6 51.2
mDPR-FT 62.7 72.5 68.4 - 48.8 56.5 59.3 71.4 58.9 51.6 49.6 64.2 59.0 59.7 68.5 80.4 69.5 - 65.0
Ours 62.3 75.3 75.0 50.4 53.8 57.0 56.3 74.0 54.1 60.0 54.7 63.2 67.1 64.3 49.9 74.2 77.2 62.3 52.3
Table 3: Comparison of nDCG@10 scores for BM25, mDPR-ZeroShot (ZS), mDPR-FineTuned (FT), and Jina-ColBERT-v2 models on the MIRACL dev set across various languages.
Table 3 compares Jina-ColBERT-v2 to BM25, mDPR, and BGE-M3. While we handily outperform BM25 and zero-shot mDPR (Zhang et al., 2023b) as expected, our model is slightly outperformed by the finetuned mDPR (Zhang et al., 2023a). For context, each mDPR-FT is only tuned on one language, rather than many like ours which may suffer to some extent from the curse of multilinguality.
mMARCO avg ar de nl es fr hi id it ja pt ru vi zh
BM-25 13.9 11.1 13.6 14.0 15.8 15.5 13.4 14.9 15.3 14.1 15.2 12.4 13.6 11.6
ColBERT-XM 25.4 19.5 27.0 27.5 28.5 26.9 23.8 26.3 26.5 24.1 27.6 25.1 22.6 24.6
Ours 31.3 27.2 33.1 33.0 34.1 33.5 30.9 31.9 33.7 27.6 33.7 29.8 28.7 30.2
Table 4: Comparison of mRR@10 scores between BM25, ColBERT-XM and Jina-ColBERT-v2 models on the mMARCO dev set across various languages.
Finally, comparing against ColBERT-XM’s zero-shot evaluation on mMARCO in Table 4, we see a strong improvement across the board, including on languages whose mMARCO training set does not occur in our pair or triplet training data (dt, hi, id, it, pt, vi).
8 Ablation Studies
In this section we present short ablation studies on modifications to three various aspects of ColBERT modeling and training.
8.1 Efficient Evaluation
Due to the compute and time costs of indexing corpora containing tens of millions of documents, evaluating every model checkpoint and ablation on every task is not feasible. Therefore, we follow recent works (Clavié, 2024; Merrick et al., 2024) by comparing models’ quality on smaller sampled-corpus versions of HotpotQA, NQ, MS MARCO, and MIRACL (Chinese, French, German, Japanese, Spanish). These sampled corpora are constructed by combining the top 250 BM25-retrieved [^2] We use the standard pre-built Lucene indices in Pyserini (Lin et al., 2021) for MIRACL found at https://github.com/castorini/pyserini, and use BM25s (Lù, 2024) for BEIR. passages with all judged passages. We observe good agreement between the sampled-corpus evaluation scores and the full-fidelity ones when used to make binary or ranking-based model comparisons, but we leave a more rigorous analysis of this observation to future work. We only use the sampled corpora for ablation studies. For the final model, we evaluate on the full version of every dataset.
8.2 Task Instructions
Inspired by the use of instruction prefixes in single-vector works like Su et al. (2022), we experimented with adding task-specific natural language instructions for retrieval (RET), and question answering (QA), and semantic text similarity (STS). However, results in Table 5 show a generally negative effect across most BEIR datasets. We hypothesize that this is because instructions are not well-suited for late interaction models, which operate at the token level. Any embedding conditioning that the instructions might provide likely becomes less effective when aggregated at the token similarity level. Furthermore, these instructions occupy valuable space within the system’s fixed token capacity.
RET QA STS
nf tc sf to db fe cf ms* fq hp* nq* ar qu sd
Mark. 32.4 59.3 67.9 19.3 35.3 67.1 18.3 34.4 37.5 25.9 40.8 37.5 86.1 18.4
Inst. 32.9 63.2 67.5 18.8 33.9 64.4 16.7 34.0 37.1 24.9 42.9 34.2 86.0 17.9
Table 5: nDCG@10 scores on BEIR datasets, grouped by task type (retrieval, question answering, and semantic text similarity) when using natural language instructions versus query/document marker tokens (default). Datasets marked with a * use the BM25-sampled corpus technique discussed in Section 8.1.
8.3 Score Normalization
BEIR MIRACL
tc hp nq ms de es fr ja zh
Baseline 78.7 36.6 58.0 45.4 57.3 40.6 50.7 63.4 63.2
+ Score Norm. 80.1 36.4 56.6 45.1 57.7 39.3 51.3 61.8 62.5
Table 6: nDCG@10 scores with and without score normalization on a retrieval-oriented subset of BEIR and MIRACL tasks. Results are performed on the BM25-sampled versions of all datasets presented except TREC-COVID (tc).
Recently, Clavié (2024) applied min-max normalization to both the student and teacher scores before computing the KL loss. This adjustment brings the score distributions of the ColBERT model and its CE teacher into closer alignment, as the original score distribution for ColBERT theoretically ranges from zero to the number of query tokens, and is model-dependent for the teacher CE. Our experiment presented in Table 6, however, shows this method to have inconclusive benefit to nDCG@10 on the BEIR and MIRACL datasets when applied to our model. We consider this result to be understandable given Clavié (2024)’s very small observed effect.
8.4 Query Augmentation Attention
BEIR MIRACL
tc hp nq ms de es fr ja zh
Baseline 77.2 70.4 54.6 37.6 33.3 40.3 35.9 54.9 34.4
+ [MASK] attn. 80.2 71.5 58.8 44.3 45.6 49.8 44.8 58.8 52.9
Table 7: nDCG@10 scores with and without query augmentation [MASK] token attention on a retrieval-oriented subset of BEIR and MIRACL tasks. Results report full-fidelity scores.
An important feature of ColBERT’s implementation is its query augmentation mechanism. By padding queries with [MASK] tokens to a uniform length, ColBERT uses BERT’s masked language modeling ability to produce additional soft term embeddings which interact with document token embeddings during MaxSim scoring. However, prior ColBERT models do not modify the attention mask to allow query tokens to attend to the mask tokens, which some hypothesize might harm generalization by making this augmentation feature too integral to the embedding process. Our controlled triplet training experiment in Table 7, however, demonstrates a positive effect across a variety of tasks, with particular benefit to non-English tasks in MIRACL. We therefore allow this attention in our training and inference.
9 Conclusion
This work presents Jina-ColBERT-v2, a capable multilingual ColBERT model that is the result of improvements to its architecture and training process. We implement modifications to the model architecture that yield efficiency gains with effectively no downside, and subsequently train it on a heterogeneous mix of data of varying tasks, languages, and supervision structures in order to bolster its performance as a general purpose retriever. Our ablation experiments demonstrate the sensitivity of ColBERT to modifications to its representations.
We hope that our work will support future multilingual ColBERT development, and prompt further exploration into the properties and optimal configuration of its query augmentation mechanism. We are also encouraged by the many inference-only optimization works on ColBERT representations, and suggest further effort be invested in tying these methods more closely with the models training objective.
10 Acknowledgement
We thank Qi Liu and Jiaxin Mao from Renmin University of China for the contributions to Jina-ColBERT-v1 and offer the insights about MRL over MRL-E for ColBERT models.